# Three results on twisted $G$-codes and skew twisted $G$-codes

### v3

@@ 2 Preliminary results
where $Aut(B)$ is the automorphisms group of $B$ and $U(B)$ denotes the subgroup given by the units of $B$, and, if we write $\prescript{g}{}{{b}}=\sigma(g)(b)$, then the following identities must be held:
@@ 3 Code Checkable
Let $G$ be a group, let $p$ be a prime number, and $\mathbb{F}_{p^{n}}$ be the finite field of $p^{n}$ elements, and let $(G,\mathbb{F}_{p^{n}},\sigma,\alpha)$ be a crossed system. Then, we know that $Aut(\mathbb{F}_{p^{n}})=Gal(\mathbb{F}_{p},\mathbb{F}_{p^{n}})\simeq C_{n}$ the cyclic group of order $n$, generated by the Frobenious map. In particular, we have that $\mathbb{F}_{p}$ is fixed by the action of $G$ and $\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ is an algebra over it prime field.
@@
It is close under this operation as $\mathbb{F}_{p^{n}}$ is an integral domain, and the associativity comes from the associativity of the product in a twisted skew group ring.
@@
The epimorphism $\psi:\mathbb{F}_{p}\hat{G}\longrightarrow\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ given by $\psi(a\overline{(b,c)})=ab\overline{c}$ for all $a\in\mathbb{F}_{p}$, $(b,c)\in\hat{G}$ and extended by linearity. It respect addition by definition so we only have to prove that it respect product, and to do so we can focus on the basis elements
@@
In [1] a question regarding where a twisted-skew G-code is code checkable is proposed.
@@
Suppose that $G$ is $p$-nilpotent, then all elements of order not a power of $p$ form a subgroup $H\leq G$. Then $\hat{H}=\mathbb{F}_{p^{n}}^{*}\times H$ is the set of all elements of order not a power of $p$ of $\hat{G}$. First, $(a,b)\in\hat{G}$ is of order $p^{c}$ if and only if $(1,1)=(a,b)^{p^{c}}=(a^{\prime},b^{p^{c}})$ with $a^{\prime}\in\mathbb{F}_{p^{n}}^{*}$, so $b$ is of order a power of $p$ and therefore $b\not\in H$ and $(a,b)\not\in\hat{H}$. In addition, if $(a,b)$ is of order $k$ and $b$ is of order $p^{c}$, then by the same argument $p^{c}|k$ and $(a^{\prime},1)=(a,b)^{p^{c}}$. Therefore we have that $k=p^{c}ord((a^{\prime},1))$. Note that $\overline{\mathbb{F}_{p^{n}}^{*}}=\mathbb{F}_{p^{n}}^{*}\times\{1\}\simeq\mathbb{F}_{p^{n}}^{*}$, so $ord((a^{\prime},1))=ord_{\mathbb{F}_{p^{n}}^{*}}(a^{\prime})|p^{n}-1$ and it is well known that $gcd(p,ord((a^{\prime},1))=1$ and $k$ is not a power of $p$
We have shown that $\hat{H}=\mathbb{F}_{p^{n}}^{*}\times H$ is the set of all elements of order not a power of $p$ of $\hat{G}$. The fact that it is a subgroup is proven analogously to 1 using the fact that $H$ is a subgroup.
@@
Let $\pi:\hat{G}\longrightarrow G$ the projection on the second component. It is clearly a epimorphism of groups. Let $\hat{H}$ be the group form by all elements of order not a power of $p$. We will show that $\pi(\hat{H})=H\leq G$ is a group form by all elements of order and not a power of $p$ in $G$. First of all, as $\hat{H}$ is a subgroup, $H$ is also a group, so we only have to prove the $p$-nilpotent part. If $a\in G$ has order not a power of $p$, then $(1,a)$ has order not a power of $p$, so $a=\pi((1,a))\in\pi(\hat{H})=H$. Now, if $a\in H$, then it is the image of $(1,a)\in\hat{H}$, and $ord(a)=ord((1,a))$ so it has order not a power of $p$. ∎
@@ 4 Abelian ideals in twisted group algebra
As $D$ is commutative, therefore $G^{\prime}$ acts by scalar. To see this, note that $x=\overline{g}\cdot\overline{h}\cdot\overline{g^{-1}}\cdot\overline{h^{-1}}$ satisfy $\psi(x)=\psi(g)\psi(h)\psi(g^{-1})\psi(h^{-1})=\psi(g)\psi(h)\psi(g)^{-1}\psi(h)^{-1}=Id$, and
@@
If $\beta_{3}\not=0$ for all $g$ then $\alpha$ is a 22-coboundary and $K^{\alpha}G\simeq KG$, so it is true as in the the group ring case it is valid. Now, if $\beta_{3}(g)=0$ for some $g$, then as $\alpha(g,h)\beta_{3}(gh)=\beta_{3}(g)\beta_{3}(h)=0$ and $\alpha(g,h)\not=0$ we have that $\beta_{3}(gh)=0$ for all $h$, so $\beta_{3}(h)=0$ for all $h\in G$, so $1v_{3}\in<v_{1},v_{2}>=M$ absurd as they are linearly independent ∎
@@ 6 Conclusion
We have solved an open question about when a twisted skew $G$-code is code checkeable. We have also generalized several results about group ring to the case of twisted group ring, such as that all twisted $G$-codes of dimension up to 3 are equivalent to an abelian group code, and also generalize a bound on the dimension and the minimum distance of a twisted $G$-code.

### v4

@@ 2 Preliminary results
where $Aut(B)$ is the group of automorphisms of $B$ and $U(B)$ denotes the subgroup given by the units of $B$, and, if we write $\prescript{g}{}{{b}}=\sigma(g)(b)$, then the following identities must hold:
@@ 3 Code Checkable
Let $G$ be a group, let $p$ be a prime number, and $\mathbb{F}_{p^{n}}$ be the finite field of $p^{n}$ elements, and let $(G,\mathbb{F}_{p^{n}},\sigma,\alpha)$ be a crossed system. Then, we know that $Aut(\mathbb{F}_{p^{n}})=Gal(\mathbb{F}_{p},\mathbb{F}_{p^{n}})\simeq C_{n}$ the cyclic group of order $n$, generated by the Frobenious map. In particular, we have that $\mathbb{F}_{p}$ is fixed by the action of $G$ and $\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ is an algebra over its prime field.
@@
It is close under the operation as $\mathbb{F}_{p^{n}}$ is an integral domain, and the associativity comes from the associativity of the product in a twisted skew group ring.
@@
The epimorphism $\psi:\mathbb{F}_{p}\hat{G}\longrightarrow\mathbb{F}_{p^{n}}^{\alpha}[G,\sigma]$ given by $\psi(a\overline{(b,c)})=ab\overline{c}$ for all $a\in\mathbb{F}_{p}$, $(b,c)\in\hat{G}$ and extended by linearity. It respects addition by definition, so we only have to prove that it respects product, and to do so we can focus on the basis elements
@@
In [1] a question regarding when twisted-skew G-code are code checkable is proposed.
@@
Suppose that $G$ is $p$-nilpotent, then all elements of order coprime with $p$ form a subgroup $H\leq G$. Then the set $\hat{H}=\mathbb{F}_{p^{n}}^{*}\times H$ is precisely the set of all elements whose order is coprime $p$ of $\hat{G}$.
First, let $(a,b)\in\hat{G}$ of order a multiple of $p$ i.e. $(a,b)=p^{c}h$ with $gcd(h,p)=1$. Then $(a,b)^{h}=(a^{\prime},b^{h})$ is of order $p^{c}$, so $b^{h}$ is of order a power of $p$, and $b\not\in H$.
In addition, if $(a,b)\in\hat{G}$ is of order coprime with $p$, by the same argument $b$ is of order coprime with $p$ and $(a,b)\in\hat{H}$.
We have shown that $\hat{H}=\mathbb{F}_{p^{n}}^{*}\times H$ is the set of all elements of order coprime with $p$ of $\hat{G}$. The fact that it is a subgroup is proven analogously to 1 using the fact that $H$ is a subgroup.
@@
Let $\pi:\hat{G}\longrightarrow G$ the projection on the second component. It is clearly a epimorphism of groups. Let $\hat{H}$ be the group form by all elements of order coprime with $p$. We will show that $\pi(\hat{H})=H\leq G$ is a group form by all elements whose order coprime with $p$ in $G$.
First of all, as $\hat{H}$ is a subgroup, $H$ is also a group, so we only have to prove the $p$-nilpotent part. If $a\in G$ has order corpime with $p$, then $(1,a)$ has order coprime with $p$, so $a=\pi((1,a))\in\pi(\hat{H})=H$. Now, if $a\in H$, then it is the image of $(1,a)\in\hat{H}$, and $ord(a)=ord((1,a))$ so it has order coprime with $p$. ∎
@@ 4 Abelian ideals in twisted group algebra
First of all, recall some basic definitions of coding theory.
Definición 4.
Let $C_{1},C_{2}\subset\mathbb{F}_{q}^{m}$ codes are said to be equivalent if there exist a diagonal matrix $D=diag(\lambda_{1},\cdots,\lambda_{m})\in Mat_{m}(\mathbb{F}_{q})$ and a permutation matrix $P$ such that $C_{1}=\{PDx;x\in C_{2}\}$. They are said to be permutation equivalent if $D=Id_{m}$.
@@
As $D$ is commutative, $G^{\prime}$ acts by scalar. To see this, note that $x=\overline{g}\cdot\overline{h}\cdot\overline{g^{-1}}\cdot\overline{h^{-1}}$ satisfy $\psi(x)=\psi(g)\psi(h)\psi(g^{-1})\psi(h^{-1})=\psi(g)\psi(h)\psi(g)^{-1}\psi(h)^{-1}=Id$, and
@@
If $\beta_{3}\not=0$ for all $g$ then $\alpha$ is a 22-coboundary and $K^{\alpha}G\simeq KG$ so the statement is true. Now, if $\beta_{3}(g)=0$ for some $g$, then as $\alpha(g,h)\beta_{3}(gh)=\beta_{3}(g)\beta_{3}(h)=0$ and $\alpha(g,h)\not=0$ we have that $\beta_{3}(gh)=0$ for all $h$, so $\beta_{3}(h)=0$ for all $h\in G$, so $1v_{3}\in<v_{1},v_{2}>=M$ absurd as they are linearly independent ∎
@@ 6 Conclusion
We have solved an open question concerning when a twisted skew $G$-code is code checkable. We have also generalized several results about group ring to the case of twisted group ring, such as that all twisted $G$-codes of dimension up to 3 are equivalent to an abelian group code, and also generalize a bound on the dimension and the minimum distance of a twisted $G$-code.